Find the values of x and y that solve the following system of equations:

5x-6y=-1 and -3x-2y=23

Answer by mananth(16946)

1. 5x-6y=-1
2. -3x-2y=23
..
multiply eq2 by 3 and subtract eq1 to eliminate y
-9x-6y-(5x-6y)= 69- (-1)
-9x-6y-5x+6y=70
-14x=70
x=70/-14
x=-5
plug the value of x n eq1
5*(-5)-6y=-1
-25-6y=-1
-6y=24
y=-4